# On-Chain Analysis of Smart Contract Dependency Risks on Ethereum

### v1

@@ IV-D1 Uniswap case study
Our goal is to examine the external interactions initiated by these contracts on-chain. Specifically, we identify and analyze all outbound calls made by documented Uniswap contracts. These contracts have interacted with a total of 4,665,557 other external addresses, we analyse the top 5 most-called target contract addresses for each documented contract, resulting in 69 distinct target as presented in Figure 6.
From the 38 distinct contracts that are not listed in the official documentation, we identify 17/38 which are in-protocol contracts created by Uniswap factory contracts or deployed by known Uniswap EOAs. These include liquidity pools, order reactors, and utility contracts. We claim that those contracts should be transparently documented in the official documentation.
We identify 21 out-of-protocol contracts, mostly callbacks, with no deployment link to Uniswap. Of the 21 out-of-protocol contracts, 5 are well-known token contracts (e.g., USDC, USDT), which are expected in a decentralized exchange context. The remaining 15 are callback contracts—addresses passed in by users during execution.
All official and in-protocol Uniswap contracts have verified source code, providing strong transparency and auditability.
Although Uniswap maintains a robust documentation portal, our analysis reveals that its token pool contracts remain undocumented despite being deployed, verified and actively called by core components. These contracts form part of the protocol’s functional surface area and should be included in the documentation to improve transparency and support safe integration.
@@ IV-D2 Lido case study
First, we extract 38 contracts from the official documentation [10] including Core Protocol, Oracle, and DAO contracts. Next, we examined the calls following Block #17244985, corresponding to the upgrade of Lido to its V2 version [8]. We compute and visualize all interactions from those core contracts. Figure 7 shows the result. It reads like Figure 6.
Among the Lido contracts, some contracts are more interconnected than others, in particular Lido DAO, stETH, and the Staking Router, as can be seen on the right hand side of Figure 7. Regarding transparency, 42 contracts in the network are included in the official documentation, whereas the other 42 are not documented. We manually analyze all of them: 15 are out-of-protocol contracts and 27 by Lido-related deployer contract or known EOAs.
Out of 27 contracts confirmed as Lido-related, 13 of them are implementation contracts, 5 are proxy contracts, 1 is a token contract and 8 contracts serve as functions such as AllowedRecipientsRegistry [26] and OracleReportSanityChecker [25]. This analysis demonstrates that the official Lido documentation is incomplete, with 27 contracts missing. We argue that all contracts which play a key role in the protocol should be documented as such.
[style=theoremstyle,frametitle=Answer to RQ4,linecolor=lightgray] The analysis of Uniswap and Lido protocols reveals significant transparency gaps in DeFi documentation. Both ecosystems show concerning discrepancies between official documentation and onchain data. 17 out of 48 and 27 out of 69 dependencies were not documented by Uniswap and Lido, respectively. For protocol developers, this underscores the urgent need to improve documentation practices and verify all contracts. For protocol users, it emphasizes the importance of understanding that official documentation may not capture the full scope of on-chain interactions and risks their funds are exposed to.
@@ V Threats to Validity
Internal Threats The blockchain analytics platforms have data inconsistencies. We tested BigQuery, Dune, and Allium and compared each platform’s data against an Erigon node. Specifically, we assessed two metrics: the total number of traces recorded in 2024 and the trace count over a segment of 1001 blocks. For the total trace count, the results we have gotten from four data sources are different. Dune and BigQuery deviated by 7.2% relative to the Erigon data, while Allium showed a negligible deviation of . Regarding the 1001-block segment, although Dune and BigQuery’s results matched those from the Erigon node, they omitted interactions with precompiled contracts; in contrast, Allium includes these interactions and the results are close to the results we have gotten from the Erigon node.
@@
External Threats Our qualitative analysis of smart contract transparency is based on two case studies: Uniswap and Lido. These findings may not generalize to all decentralized applications. To mitigate this risk, we deliberately selected two of the most widely used and highly audited DeFi protocols. Uniswap is recognized as the top decentralized exchange protocol, and Lido is the foremost liquid staking protocol on Ethereum [2]. Our analysis spans 153 contracts associated with these platforms, offering a robust foundation for understanding the transparency challenges faced by prominent DApps.
@@ VI Related Work
Graph Analysis. Khan’s survey [34] maps the related work on graph analysis of Ethereum, up to 2021. Early notable papers include [51] and [38] which look at graph theoretic metrics such as in and out-degree in the Ethereum transfer graph. Lyu et al. [40] used contract dependency graph with automated analysis to analyze the impact of vulnerable contract. Beyond aggregated metrics, Ao et al. [16] made a deep dive on graph metrics for the Aave protocol. Lee et al. [36] study different interaction types on the Ethereum incl. contract-to-contract interactions. None of this related work analyses smart contract dependencies as we do in this paper.
Smart Contract Dependencies. A couple of tools has been developed to assist smart contract developers by providing graph visualizations of execution flows between smart contracts namely Smart-Graph [44] and MindTheDapp [31]. They focus on analyzing individual DApps instead of the entire smart contract ecosystem within a blockchain (what this work does). Aufiero et al. [18] used static analysis to study function calls within DApps and found that most DApp functionality is concentrated in a small subset of functions, particularly those in token contracts. Several studies analyze on-chain data to map smart contract compositions. Kitzler et al. [35] used on-chain transaction data to conducted a topology analysis for DeFi protocols in the time span from January 2021 to August 2021. Fröwis and Böhme [28] measure the apparent trustless nature of smart contracts through a call graph of dependencies, finding out that two out of five smart contracts require trust in at least one-third party. Ferreti and D’Angelo [27] analyze the interactions between smart contracts through network modeling using historical transactions. Serena et al. [48] analyzed the transaction record of four DLTs and specific characteristics of a network. Harrigan et al. [29] study token compositions by constructing a graph of tokens with the events emitted by them. To our knowledge, we are the first to provide a large-scale, up-to-date analysis of smart contract dependencies across the entire Ethereum blockchain, offering novel insights into the ecosystem’s scale and transparency, as demonstrated in RQ1 and RQ4.
Smart Contract Upgradability. Huang et al. [30] used bytecode-based method to examine the prevalence of upgradeable smart contracts on Ethereum. Qasse et al. [45] examined the likeness of an upgradeable contract to be upgraded. Bodell III et al. [32] implemented static analysis framework to analyzed proxy-based upgradeable smart contract. Li et al. [37] used bytecode and transaction information to identify upgradeable contracts. Ruaro et al. [46] analyzed storage collision and identified potentially vulnerable contracts. None of them studied the fundamental risks of upgradable popular contracts, as we have done in RQ3.
@@ Acknowledgements
This work was partially supported by the Swedish Foundation for Strategic Research (SSF); the Ethereum Foundation; the Swedish Governmental Agency for Innovation Systems (Vinnova). We thank Christof Torres for valuable feedback on the work.

### v2

@@ IV-D1 Uniswap case study
Our goal is to examine the external interactions initiated by these contracts on-chain. Specifically, we identify and analyze all outbound calls made by documented Uniswap contracts.
In total, we gathered 47 documented Uniswap addresses, of which 34 have interactions. These contracts have interacted with 4,701,217 distinct external addresses. To streamline our analysis, we examine the top 5 most frequently called target contract addresses for each documented contract, resulting in a set of 77 unique addresses, as illustrated in Figure 6.
From the 43 distinct contracts that are not listed in the official documentation, we identify 18 in-protocol contracts created by Uniswap factory contracts or deployed by known Uniswap EOAs. These include liquidity pools and utility contracts. We claim that these contracts should be transparently documented in the official documentation.
We identify 25 out-of-protocol contracts, mostly callbacks, with no deployment link to Uniswap. Of the 25 out-of-protocol contracts, 7 are well-known token contracts (e.g., USDC, USDT), which are expected in a decentralized exchange context. The remaining 18 are callback contracts, addresses passed in by users during execution.
All official and in-protocol Uniswap contracts have verified source code, providing strong source code transparency and auditability.
Although Uniswap maintains a robust documentation portal, our analysis reveals that its token pool contracts remain undocumented despite being deployed, verified, and actively called by core components. These contracts form part of the protocol’s functional surface area, and their addresses should be included in the documentation to improve transparency and support safe integration.
@@ IV-D2 Lido case study
First, we extract 39 contracts from the official documentation [10] including Core Protocol, Oracle, and DAO contracts. Next, we examined the calls following Block #17244985, corresponding to the upgrade of Lido to its V2 version [8]. We compute and visualize all interactions from those core contracts. Figure 7 shows the result. It reads like Figure 6.
Among the Lido contracts, some contracts are more interconnected than others, in particular Lido DAO, stETH, and the Staking Router, as can be seen on the right-hand side of Figure 7. Regarding transparency, 53 contracts in the network are included in the official documentation, whereas the other 41 are not documented. We manually analyze all of them: 16 are out-of-protocol contracts and 25 are by Lido-related deployer contracts or known EOAs.
Out of 25 contracts confirmed as Lido-related, 16 of them are implementation contracts, 4 are proxy contracts, 1 is a token contract, and 4 serve as functions such as OracleReportSanityChecker [25]. This analysis demonstrates that the official Lido documentation is incomplete, with 25 contracts missing. We argue that all contracts which play a key role in the protocol should be documented as such.
[style=theoremstyle,frametitle=Answer to RQ4,linecolor=lightgray] The analysis of Uniswap and Lido protocols reveals significant transparency gaps in DeFi documentation. Both ecosystems show concerning discrepancies between official documentation and on-chain data. 18 out of 52 and 25 out of 78 dependencies were not documented by Uniswap and Lido, respectively. For protocol developers, this underscores the urgent need to improve deployment documentation practices. For protocol users, it emphasizes the importance of understanding that official documentation may not capture the full scope of on-chain interactions and risks their funds are exposed to.
@@ V Threats to Validity
Internal Threats. The blockchain analytics platforms have data inconsistencies. We tested BigQuery, Dune, and Allium and compared each platform’s data against an Erigon node. Specifically, we assessed two metrics: the total number of traces recorded in 2024 and the trace count over a segment of 1001 blocks. For the total trace count, the results we have gotten from four data sources are different. Dune and BigQuery deviated by 7.2% relative to the Erigon data, while Allium showed a negligible deviation of . Regarding the 1001-block segment, although Dune and BigQuery’s results matched those from the Erigon node, they omitted interactions with precompiled contracts; in contrast, Allium includes these interactions and the results are close to the results we have gotten from the Erigon node.
@@
External Threats. Our qualitative analysis of smart contract transparency is based on two case studies: Uniswap and Lido. These findings may not generalize to all decentralized applications. To mitigate this risk, we deliberately selected two of the most widely used and highly audited DeFi protocols. Uniswap is recognized as the top decentralized exchange protocol, and Lido is the foremost liquid staking protocol on Ethereum [2]. Our analysis spans 153 contracts associated with these platforms, offering a robust foundation for understanding the transparency challenges faced by prominent DApps.
@@ VI Related Work
Graph Analysis. Khan’s survey [33] maps the related work on the graph analysis of Ethereum, up to 2021. Early notable papers include [50] and [37] which look at graph theoretic metrics such as in and out-degree in the Ethereum transfer graph. Lyu et al. [39] used contract dependency graph with automated analysis to analyze the impact of vulnerable contracts. Beyond aggregated metrics, Ao et al. [16] made a deep dive on graph metrics for the Aave protocol. Lee et al. [35] study different interaction types on the Ethereum incl. contract-to-contract interactions. None of this related work analyzes smart contract dependencies as we do in this paper.
Smart Contract Dependencies. A couple of tools have been developed to assist smart contract developers by providing graph visualizations of execution flows between smart contracts namely Smart-Graph [43] and MindTheDapp [30]. They focus on analyzing individual DApps instead of the entire smart contract ecosystem within a blockchain (what this work does). Aufiero et al. [18] used static analysis to study function calls within DApps and found that most DApp functionality is concentrated in a small subset of functions, particularly those in token contracts. Several studies analyze on-chain data to map smart contract compositions. Kitzler et al. [34] used on-chain transaction data to conducted a topology analysis for DeFi protocols in the time span from January 2021 to August 2021. Fröwis and Böhme [27] measure the apparent trustless nature of smart contracts through a call graph of dependencies, finding out that two out of five smart contracts require trust in at least one-third party. Ferreti and D’Angelo [26] analyze the interactions between smart contracts through network modeling using historical transactions. Serena et al. [47] analyzed the transaction record of four DLTs and specific characteristics of a network. Harrigan et al. [28] study token compositions by constructing a graph of tokens with the events emitted by them. To our knowledge, we are the first to provide a large-scale, up-to-date analysis of smart contract dependencies across the entire Ethereum blockchain, offering novel insights into the ecosystem’s scale and transparency, as demonstrated in RQ1 and RQ4.
Smart Contract Upgradability. Huang et al. [29] used a bytecode-based method to examine the prevalence of upgradeable smart contracts on Ethereum. Qasse et al. [44] examined the likeness of an upgradeable contract to be upgraded. Bodell III et al. [31] implemented a static analysis framework to analyzed proxy-based upgradeable smart contract. Li et al. [36] used bytecode and transaction information to identify upgradeable contracts. Ruaro et al. [45] analyzed storage collision and identified potentially vulnerable contracts. None of them studied the fundamental risks of upgradable popular contracts, as we have done in RQ3.
@@ Acknowledgements
This work was partially supported by the Swedish Foundation for Strategic Research (SSF); the Ethereum Foundation; the Swedish Governmental Agency for Innovation Systems (Vinnova). We thank Christof Torres for the valuable feedback on the work.